# Towards Competent AI for Fundamental Analysis in Finance: A Benchmark Dataset and Evaluation

### v1

@@ 4.1 Table Assessment
    $D_{\theta}(p,t)=\begin{cases}\frac{||p-t||}{||t||},&\text{if }\frac{||p-t||}{||t||}\leq\theta\\ 1,&\text{otherwise}\end{cases}$ (3)
@@ 4.2 Reasoning Assessment
  • Accuracy: Are facts correctly interpreted?
  • Depth of analysis: Are findings linked together into an integrated financial diagnosis?
  • Financial insight: Are interpretations thoughtful and informative?
@@ 5.1 Experimental Setup
Experiment Settings For open-sourced LLMs, we evaluate them through NVIDIA NIM API. The models are used with default parameter settings, and maxtokens is set to the maximum value to prevent output truncation during inference. Since LLMs can not handle PDF input directly, we use PyMuPDF  [^2] https://pymupdf.readthedocs.io/en/latest/ as the PDF extractor, which outperforms other alternatives according to our experiment, see Appendix A. During the evaluation, we set the text matching threshold $\tau=1$, ensuring that all key pairs are considered eligible during the assignment phase. We set the numerical error threshold $\theta=0$, reflecting high precision requirements in finance.
@@ 5.2 Main Results
Table 1 shows the main results of baseline LLMs on the test set of FinAR-Bench. In the information extraction task, large-sized LLMs achieve near-perfect scores, while medium-sized models remain competent. In contrast, small-sized LLMs exhibit significant performance degradation, primarily due to their limited capacity to process and generate long texts. In the indicator computation task, all models regardless of size perform poorly, reflecting a general deficiency in precise numerical computation. To further investigate LLMs’ capability in approximate numerical computation, we perform additional evaluation by varying error tolerance threshold (refer to Table 3). Performance on the logic reasoning task surpasses that of the indicator computation task, largely due to a greater tolerance for numerical imprecision. This finding aligns with patterns observed in LLM-generated financial analysis reports, where the logical reasoning often appears sound even when the quantitative details are inaccurate. These results suggest that LLMs may still offer valuable insights in financial contexts, despite their limitations in exact calculation. Moreover, model performance is consistently lower in the PDF setting compared to the text setting as the variability of PDF layout adds complexity.
Table 1: Precision and recall for models across three tasks under PDF and text inputs. Missing values are due to the model’s context length limitation or its inability to generate long content.
@@
PDF Text PDF Text PDF Text
P R P R P R P R P R P R
L GPT-4o 94.90 93.89 98.39 98.39 33.78 33.78 34.38 34.38 60.61 60.26 63.23 63.23
GPT-o1 96.08 96.08 100.00 100.00 33.78 33.78 34.76 34.76 85.73 85.73 86.20 86.20
DeepSeek-v3 93.42 93.42 99.98 99.98 35.28 35.28 38.26 38.26 58.35 58.72 63.11 63.11
DeepSeek-r1 96.44 96.20 100.00 99.93 48.09 47.63 49.31 49.31 84.40 73.40 83.83 76.28
Llama-3.1-405b-instruct 95.67 92.61 99.03 99.03 18.65 18.63 20.83 20.80 61.78 55.40 64.90 63.84
Mixtral-8*22b-instruct-v0.1 85.82 85.03 97.34 98.19 12.24 12.22 21.96 21.94 54.25 54.25 48.05 47.94
@@
We conduct a tournament-style evaluation of reasoning outputs using Doubao-1.5-thinking-pro as the judge. Only large-sized models with recall scores above 60% are included in the competition. To avoid order bias, each pair of models competes in two matches: one where model A is evaluated against model B, and another with the order reversed (model B v.s. model A). This setup yields a total of 1800 competitions among five LLMs. As shown in Table 2, GPT-o1 and DeepSeek-r1 leads the ranking with 647 and 603 wins, respectively. GPT-4o shows moderate performance, while Llama-3.1-405b-instruct lags far behind.
Table 2: Tournament ranking for the reasoning analysis.
Model Score Rank Avg. Generated Tokens per Analysis
GPT-o1 647 1 9681.86
DeepSeek-r1 603 2 5400.77
DeepSeek-v3 316 3 2491.16
GPT-4o 218 4 2024.49
Llama-3.1-405b-instruct 16 5 1741.11
@@ 5.3.3 Effect of Knowledge Augmentation
The financial knowledge encoded in LLMs may not align with the standard formulas we use to calculate financial indicators. We investigate whether explicitly providing calculation equations improves LLMs’ performance on financial indicator tasks. In the "enhanced prompt" setting, we include exact calculation equations (e.g., "Net Profit Margin = Net Profit / Revenue"), while in the "basic prompt" setting, models are asked to compute indicators without supplementary information. Figure 2 reports the evaluation under different numerical error threshold $\theta$. Most medium-sized and large-sized models exhibit substantial improvements in recall when provided with enhanced prompts. In particular, under the numerical error threshold $\theta=0.01$ setting, GPT-o1 improves from 79.76% to 98.00%, and DeepSeek-r1 from 76.43% to 94.00%, demonstrating the effectiveness of knowledge augmentation in improving numerical performance. In contrast, smaller models (e.g., Mistral-7b, DeepSeek-r1-distill-llama-8b) show minimal or even negative gains. This may stem from limited capacity in handling long structured inputs, where increased prompt complexity leads to confusion and degraded numerical reasoning. Notably, models with stronger reasoning abilities—such as DeepSeek-r1, GPT-o1, and Qwq-32b—benefit the most from enhanced prompts. These models can better internalize explicit formula knowledge, leading to evidently improved performance.
@@ 5.4 Case Study
To evaluate the quality of financial reasoning produced by LLMs, we conduct a case study by manually reviewing outputs of GPT-o1 and DeepSeek-r1 in the logical reasoning task for a selected company. Reviews are summarized in Table 4. We find that current models still show clear limitations in financial analysis. GPT-o1 has fewer logic errors than DeepSeek-r1 and both models perform poorly in terms of analytical depth and financial insight. This suggests that further optimization and domain-specific enhancement are needed for LLMs in financial applications.
@@
Accuracy - Interpretations are basically correct but lacks concrete numerical references.
- Partial interpretation: "net profit < operating cash flow" $\Rightarrow$ "strong cash collection capability" — Ignores other influential factors of cash collection capability.
- Logic contradiction: "net profit margin (1.96% $\rightarrow$ 4.98%) declined"; A small margin increase (7.98% $\rightarrow$ 10.98%) is interpreted as "significant deterioration".
@@
- Misunderstanding of terminology: "…, indicating a diminished capacity of cash flow to cover profits" — Cash flow does not need to cover profits. It only needs to cover loan repayments and interest.
@@
- Tends to force causal links between unrelated indicators and cause the incorrect analysis: "Inventory decreased 19.5% but still higher than fixed assets $\Rightarrow$ inventory obsolescence risk" — In fact, inventory decrease is more likely due to sales improvement.
- Incorrect causal inference: "Net financing cash flow was -1.28B due to debt repayment (245.92B) exceeding new borrowings (246.56B) $\Rightarrow$ "active deleveraging" — The simultaneous occurrence of repayments and borrowings, offsetting one another, seems to reflect refinancing, not deleveraging.
@@
Conclusion - Basic financial understanding, student level. - No practical use value.
@@ Acknowledgment
Appendix A PDF Extractor Results

### v2

@@ 4.1 Table Assessment
[…]
@@ 4.2 Reasoning Assessment
  • Consistency: Evaluate whether each report is internally coherent and logically sound. This includes checking the alignment between numerical data, identified trends, and stated conclusions, as well as the adherence to standard financial reasoning practices.
  • Depth of Analysis: Assess the extent to which each report transcends basic fact listing to construct an integrated financial diagnosis. A strong analysis should meaningfully connect multiple financial indicators and reveal the relationships among key metrics.
  • Financial Insight: Determine whether the report offers thoughtful and non-obvious interpretations. Reports with high insight demonstrate financial expertise, critical thinking, and the ability to explain underlying causes, implications, or strategic considerations behind the numbers.
@@ 5.1 Experimental Setup
Experiment Settings For open-sourced LLMs, we evaluate them through NVIDIA NIM API. The models are used with default parameter settings, and maxtokens is set to the maximum value to prevent output truncation during inference. Since LLMs can not handle PDF input directly, we use PyMuPDF  [^2] https://pymupdf.readthedocs.io/en/latest/ as the PDF extractor, which outperforms other alternatives according to our experiment, refer to Appendix D. During the evaluation, we set the text matching threshold $\tau=1$, ensuring that all key pairs are considered eligible during the assignment phase. We set the numerical error threshold $\theta=0$, reflecting high precision requirements in finance.
@@ 5.2 Main Results
Table 1 shows the main results of baseline LLMs on the test set of FinAR-Bench under the text format. Results under the PDF format are provided in Appendix D. In the information extraction task, large-sized LLMs achieve near-perfect scores, while medium-sized models remain competent. In contrast, small-sized LLMs exhibit significant performance degradation, primarily due to their limited capacity to process and generate long texts. In the indicator computation task, all models regardless of size perform poorly, reflecting a general deficiency in precise numerical computation. To further investigate LLMs’ capability in approximate numerical computation, we perform additional evaluation by varying error tolerance threshold (refer to Table 3). Performance on the logic reasoning task surpasses that of the indicator computation task, largely due to a greater tolerance for numerical imprecision. This finding aligns with patterns observed in LLM-generated financial analysis reports, where the logical reasoning often appears sound even when the quantitative details are inaccurate. These results suggest that LLMs may still offer valuable insights in financial contexts, despite their limitations in exact calculation. Moreover, model performance is consistently lower in the PDF setting compared to the text setting as the variability of PDF layout adds complexity.
Table 1: Precision and recall of all models under text input.
@@
P R P R P R
L GPT-4o 98.39 98.39 34.38 34.38 63.23 63.23
GPT-o1 100.00 100.00 34.76 34.76 86.20 86.20
DeepSeek-v3 99.98 99.98 38.26 38.26 63.11 63.11
DeepSeek-r1 100.00 99.93 49.31 49.31 83.83 76.28
Llama-3.1-405b-instruct 99.03 99.03 20.83 20.80 64.90 63.84
Mixtral-8*22b-instruct-v0.1 97.34 98.19 21.96 21.94 48.05 47.94
@@
We conduct a tournament-style evaluation of reasoning outputs using Doubao-1.5-thinking-pro, Qwen-Max and Gemini-2.5-Pro as the judges. Only large-sized models with recall scores above 60% are included in the competition. To avoid order bias, each pair of models competes in two matches: one where model A is evaluated against model B, and another with the order reversed (model B v.s. model A). This setup yields a total of 1800 competitions among 5 LLMs per judge. As shown in Table 2, GPT-o1 and DeepSeek-r1 leads the ranking with 634.3 and 624.5 wins, respectively. DeepSeek-r1 and GPT-4o shows moderate performance, while Llama-3.1-405b-instruct lags far behind.
Table 2: Scores from three LLM judges and summary statistics. Abbreviations: D = Doubao-1.5-pro, Q = Qwen-Max, G = Gemini-2.5-Pro, Avg = Average score, Tok = Avg. generated tokens, Rk = Rank.
Model D Q G Tok Avg / Rk
GPT-o1 649 637 617 9681.9 634.3 / 1
DeepSeek-r1 609 622.5 642 5400.8 624.5 / 2
DeepSeek-v3 293.5 275.5 202.5 2491.2 257.2 / 3
GPT-4o 232.5 245 223.5 2024.5 233.7 / 4
Llama-3.1-405b 16 20 115 1741.1 50.3 / 5
@@ 5.3.3 Effect of Knowledge Augmentation
The financial knowledge encoded in LLMs may not align with the standard formulas we use to calculate financial indicators. We investigate whether explicitly providing calculation equations improves LLMs’ performance on financial indicator tasks. In the "enhanced prompt" setting, we include exact calculation equations (e.g., "Net Profit Margin = Net Profit / Revenue"), while in the "basic prompt" setting, models are asked to compute indicators without supplementary information. Figure 2 illustrates the results of the evaluation under a strict numerical error threshold of $\theta=0$. Most medium-sized and large-sized models exhibit substantial improvements in recall when provided with enhanced prompts. In contrast, smaller models (e.g., Mistral-7b, DeepSeek-r1-distill-llama-8b) show minimal or even negative gains. This may stem from limited capacity in handling long structured inputs, where increased prompt complexity leads to confusion and degraded numerical reasoning. Notably, models with stronger reasoning abilities, such as DeepSeek-r1, GPT-o1, and Qwq-32b benefit the most from enhanced prompts. These models can better internalize explicit formula knowledge, leading to evidently improved performance.
@@ 5.4 Case Study
To evaluate the quality of financial reasoning produced by LLMs, we conduct a case study by manually reviewing outputs of GPT-o1 and DeepSeek-r1 in the logical reasoning task for a selected company. This evaluation is carried out by an financial expert, ensuring a comprehensive and credible assessment of the models’ outputs. Reviews are summarized in Table 4. We find that current LLMs still show clear limitations in financial analysis. GPT-o1 has fewer logic errors than DeepSeek-r1 and both models perform poorly in terms of analytical depth and financial insight. This suggests that further optimization and domain-specific enhancement are needed for LLMs in financial applications.
@@
Consistency - Interpretations are basically correct but lacks concrete numerical references.
- Partial interpretation: "net profit << operating cash flow" implies "strong cash collection capability". It ignores other influential factors of cash collection capability.
- Logic contradiction: A small margin increase (from 7.98% to 10.98%) is interpreted as "significant deterioration".
@@
- Misunderstanding of terminology: "…, indicating a diminished capacity of cash flow to cover profits". Cash flow does not need to cover profits. It only needs to cover loan repayments and interest.
@@
- Tends to force causal links between unrelated indicators and cause the incorrect analysis: "Inventory decreased 19.5% but still higher than fixed assets" is interpreted as inventory obsolescence risk. In fact, inventory decrease is more likely due to sales improvement.
- Incorrect causal inference: "Net financing cash flow was -1.28B due to debt repayment (245.92B) exceeding new borrowings (246.56B) " leading to the conclusion of "active deleveraging". The simultaneous occurrence of repayments and borrowings, offsetting one another, seems to reflect refinancing, not deleveraging.
@@
Conclusion - Basic financial understanding. - No practical use value.
@@ Acknowledgment
Appendix A Descriptive Statistics of FinAR-Bench
FinAR-Bench consists of a total of 1,170 financial analysis tasks from 100 companies, divided into a development set with 90 companies and a test set with 10 companies. Each company contributes 13 tasks: 6 factual extraction tasks, 6 indicator computation tasks, and 1 reasoning task. Factual and indicator tasks are defined over six task sizes (1, 2, 4, 8, 16, 32 tables), enabling evaluation of LLM performance under different task sizes. The reasoning task requires logical interpretation over given conditions. All tasks are constructed from real-world annual reports covering fiscal years 2022 and 2023. The development set is publicly released, while the test set is reserved for benchmarking and leaderboard evaluations.
Table 5: Summary Statistics of FinAR-Bench
Metric Value
Total Companies 100
Dev / Test Split 90 / 10
Tasks per Company 13
Total Tasks 1,170
Fact Tasks (per company) 6
Indicator Tasks (per company) 6
Reasoning Tasks (per company) 1
Fiscal Years Covered 2022, 2023
Appendix B Task Details
This appendix summarizes the financial items, indicators, and reasoning conditions used in the FinAR-Bench evaluation. It consists of three components:
Financial Items
Table 6 lists the financial items involved in all tasks.
Table 6: Financial Items
No. Item No. Item
1 Total Revenue 17 Net Value of Fixed Assets
2 Total Cost of Goods Sold 18 Construction in Progress
3 Revenue 19 Short-term Loans
4 Operating Cost 20 Long-term Debt
5 Total Profit 21 Bonds Payable
6 Total Assets 22 Net Income
7 Total Liabilities 23 Net Income Attributable to Parent
8 Total Current Assets 24 Selling Expenses
9 Total Non-Current Assets 25 Administrative Expenses
10 Total Current Liabilities 26 Financial Expenses
11 Total Long-Term Liabilities 27 Accounts Receivable
12 Total Shareholders’ Equity Including Minority Interest 28 Accounts Payable
13 Total Shareholders’ Equity Excluding Minority Interest 29 Payroll Payable
14 Money Capital 30 Net Cash Provided by Operating Activities
15 Goodwill 31 Net Cash Used in Investing Activities
16 Inventories 32 Net Cash Provided Used in Financing Activities
Financial Indicators and Calculation Formulas
Table 7 summarizes the financial indicators used in our evaluation tasks, along with their corresponding calculation formulas.
Table 7: Financial Indicators and Calculation Formulas
No. Indicator Formula
1 Return on Equity Net Income Attributable to Parent / Average Equity Attributable to Parent
2 Return on Assets Net Income / Average Total Assets
3 Gross Profit Margin (Operating Revenue - Operating Cost) / Operating Revenue
4 Net Profit Margin Net Income / Operating Revenue
5 Debt to Assets Ratio Total Liabilities / Total Assets
6 Current Ratio Current Assets / Current Liabilities
7 Quick Ratio (Current Assets - Inventories - Prepayments - Current Portion of Non-current Assets - Other Current Assets) / Current Liabilities
8 Period Expense Ratio Gross Profit Margin - Net Profit Margin
9 Equity Multiplier Total Assets / Total Shareholders’ Equity
10 Equity Ratio Total Liabilities / Equity Attributable to Parent
11 Inventory Turnover Days 360 / Inventory Turnover Ratio
12 Accounts Receivable Turnover Days 360 / Accounts Receivable Turnover Ratio
13 Accounts Payable Turnover Days 360 / Accounts Payable Turnover Ratio
14 Operating Cycle Inventory Turnover Days + Accounts Receivable Turnover Days
15 Total Asset Turnover Ratio Operating Revenue / Average Total Assets
16 Inventory Turnover Ratio Operating Cost / Average Inventories
17 Accounts Receivable Turnover Ratio Operating Revenue / Average Accounts Receivable
18 Accounts Payable Turnover Ratio (Ending Inventories + Operating Cost - Beginning Inventories) / Average Accounts Payable
19 Current Asset Turnover Ratio Operating Revenue / Average Current Assets
20 Fixed Asset Turnover Ratio Operating Revenue / Average Net Fixed Assets
21 Net Profit Growth Rate (Current Year Net Income - Previous Year Net Income) / Previous Year Net Income
22 Operating Revenue Growth Rate (Current Year Revenue - Previous Year Revenue) / Previous Year Revenue
23 Net Operating Cash Flow Growth Rate (Current Year Operating Cash Flow - Last Year Operating Cash Flow) / Last Year Operating Cash Flow
24 Accounts Receivable Growth Rate (Ending Accounts Receivable - Beginning Accounts Receivable) / Beginning Accounts Receivable
25 Net Cash Flow to Net Profit Ratio Net Operating Cash Flow / Net Income
26 Operating Cash Inflow to Revenue Ratio Cash Received from Sale of Goods and Services / Operating Revenue
27 Selling Expenses to Revenue Ratio Selling Expenses / Operating Revenue
28 Administrative Expenses to Revenue Ratio Administrative Expenses / Operating Revenue
29 Financial Expenses to Revenue Ratio Financial Expenses / Operating Revenue
30 Goodwill to Total Assets Ratio Goodwill / Total Assets
31 Accounts Receivable to Total Assets Ratio Accounts Receivable / Total Assets
32 Accounts Payable to Total Liabilities Ratio Accounts Payable / Total Liabilities
33 Inventory to Total Assets Ratio Inventories / Total Assets
34 Fixed Assets to Total Assets Ratio Net Fixed Assets / Total Assets
35 Current Assets to Total Assets Ratio Current Assets / Total Assets
36 Current Liabilities to Total Liabilities Ratio Current Liabilities / Total Liabilities
37 Non-current Assets to Total Assets Ratio Non-current Assets / Total Assets
38 Non-current Liabilities to Total Liabilities Ratio Non-current Liabilities / Total Liabilities
Reasoning Task Conditions
Table 8 lists the reasoning task conditions in our benchmark. These tasks require models to interpret trends, compare values, or apply threshold-based logic to financial indicators across different fiscal years.
Table 8: Reasoning Task Conditions
Trend-Based Judgments
1/2 Return on Equity (ROE) is increasing / decreasing
3/4 Return on Assets (ROA) is increasing / decreasing
5/6 Net Profit Margin is increasing / decreasing
7/8 Gross Profit Margin is increasing / decreasing
9/10 Operating Expense Ratio is increasing / decreasing
11/12 Equity Multiplier is increasing / decreasing
13/14 Debt to Assets Ratio is increasing / decreasing
15/16 Current Ratio is increasing / decreasing
17/18 Quick Ratio is increasing / decreasing
19/20 Administrative Expenses to Revenue Ratio is increasing / decreasing
21/22 Selling Expenses to Revenue Ratio is increasing / decreasing
23/24 Financial Expenses to Revenue Ratio is increasing / decreasing
25/26 Current Assets to Total Assets Ratio is increasing / decreasing
27/28 Non-current Assets to Total Assets Ratio is increasing / decreasing
29/30 Current Liabilities to Total Liabilities Ratio is increasing / decreasing
31/32 Long-term Liabilities to Total Liabilities Ratio is increasing / decreasing
33/34 Goodwill to Total Assets Ratio is increasing / decreasing
35/36 Accounts Receivable to Total Assets Ratio is increasing / decreasing
37/38 Accounts Payable to Total Liabilities Ratio is increasing / decreasing
39/40 Net Fixed Assets to Total Assets Ratio is increasing / decreasing
Threshold-Based Judgments
41/42 Operating Revenue Growth Rate >> 0 / << 0
43/44 Net Profit Growth Rate >> 0 / << 0
45/46 Operating Cash Flow Growth Rate >> 0 / << 0
47/48 Accounts Receivable Growth Rate >> 0 / << 0
49/50 Net Operating Cash Flow >> 0 / << 0
51/52 Net Investing Cash Flow >> 0 / << 0
53/54 Net Financing Cash Flow >> 0 / << 0
Comparative Judgments
55/56 Operating Revenue Growth Rate >> / << Net Profit Growth Rate
57/58 Net Profit Growth Rate >> / << Operating Cash Flow Growth Rate
59/60 Inventories >> / << Net Fixed Assets
61/62 Net Profit >> / << Net Operating Cash Flow
63/64 Selling Expenses >> / << Administrative Expenses
Appendix C Performance of Large Models on PDF Format
In addition to the main experiments on clean text-format inputs, we evaluate the robustness of large-scale LLMs (sizes >100B) on PDF-format annual reports, which more closely resemble real-world usage scenarios. Small and medium-sized models are excluded due to the model’s context length limitation or its inability to generate long content.
Table 9 presents the precision and recall of six large models across the three task types. While information extraction remains relatively strong (e.g., DeepSeek-r1 reaches 96.44% precision), indicator computation shows significant degradation, with most models achieving under 35% recall. Logic reasoning tasks remain highly variable, with GPT-o1 reaching 85.73% recall, while Mixtral-8x22B drops below 55%.
These results highlight the difficulty of parsing structured content directly from PDFs and the performance bottlenecks in numerical computation and reasoning when layout noise is introduced. Improving layout-aware modeling and financial pretraining remains a key direction for future work.
Table 9: Precision and recall of the largest models (Size=L) under PDF input.
P R P R P R
L GPT-4o 94.90 93.89 33.78 33.78 60.61 60.26
GPT-o1 96.08 96.08 33.78 33.78 85.73 85.73
DeepSeek-v3 93.42 93.42 35.28 35.28 58.35 58.72
DeepSeek-r1 96.44 96.20 48.09 47.63 84.40 73.40
Llama-3.1-405b-instruct 95.67 92.61 18.65 18.63 61.78 55.40
Mixtral-8*22b-instruct-v0.1 85.82 85.03 12.24 12.22 54.25 54.25
Appendix D PDF Extractor Results